OpenStudy (anonymous):

As a car passes the point A on s straight road, its speed is 10 m/s. The car moves with constant acceleration a m/s² along the road for T seconds until it reaches the point B, where its speed is V m/s. The car travels at this speed for a further 10 seconds, when it reaches point C. From C it travels for a further T seconds with constant acceleration 3a m/a² until it reaches a speed of 20 m/s at the point D. Show that V= 125. Given that the distance between A and D is 675 m, Find the values of a and T.

OpenStudy (anonymous):

1) newton's law v = u + at V = 10 + at till C no acceleration , so spped is same then at D speed is 20 , so 20 = u + at , in this case u is 10 + at 20 = 10 + at + 3at solve it & u get at

OpenStudy (anonymous):

then use that to solve the rest